OpenStudy (anonymous):

Helium is pumped into a spherical balloon at a rate of 3 cubic feet per second. How fast is the radius increasing after 2 minutes?

OpenStudy (phi):

I would start with the equation for the volume of a sphere \[ V = \frac{4}{3} \pi r^3 \] take the derivative with respect to time: \[ \frac{dV}{dt} = 4 \pi r^2 \frac{dr}{dt} \] next, you need to find the volume of the balloon after being filled for 2 minutes (120 seconds) use that to solve for r. they tell you dV/dt = 3 cu ft/sec you now have enough info to solve for dr/dt

OpenStudy (accidentalaichan):

3/1 proportion! \[\frac{ 3 }{ 1 } \times60 \neq \frac{ 2 }{ x }\] 3 x 60 then multiply that by 2

OpenStudy (accidentalaichan):

or 3 x 120

OpenStudy (accidentalaichan):

360 cubic feet is my answer